# Bruno Ferman  [^2] Sao Paulo School of Economics - FGV    Lucas Finamor [^3] Sao Paulo School of Economics - FGV

### v1

@@ 4.2 Heterogeneity
The remaining columns investigate heterogeneity by correctly clustering standard errors in the screening questions (columns 3 and 4), and above or below median score in Q1–Q2 (columns 5 and 6). Along these dimensions, the results are very similar. Groups we expect to have better performance (who know how to cluster standard errors and with higher initial scores) have higher probabilities of correctly taking the 99 missing values regardless of the results the coding error would generate. However, these advantages do not make them less vulnerable to the bias. On the contrary, the point estimates are larger for the more trained subgroups although the differences are only significant at 10% for the comparison between those who did or did not correctly clustered the standard errors). Table A.5 shows the results for the sub-sample with master’s degree or above and for those who have taken an econometrics course and we see similar results. [^12] We do not show the results for the sub-samples without master’s and without an econometrics course because they are very small. Only 38 individuals do not have a master’s degree or above and 75 have not taken an econometrics course. In addition to these results, Appendix Table A.3 shows results separately for each wave.
@@ 5 Discussion
Our findings are particularly relevant for placebo tests, where researchers typically expect to find no significant effects—and where expected results are often also seen as favorable. In such cases, if a coding error leads to a statistically insignificant placebo result, researchers may interpret this as confirmation that the test “worked” and may forgo further scrutiny. As a consequence, coding errors may lead to an excess of false-negative findings, masking potential violations of identifying assumptions even when researchers are acting in good faith.
@@ 6 Conclusion
In recent years, the Economics profession has seen increased concerns about the reproducibility and replicability of research findings. Many journals set up policies requiring data and code availability to increase research transparency. In this paper, we experimentally test whether individuals are more likely to find coding errors when they lead to non-expected results. We find that the probability of spotting a simple and frequent coding error increases by 20% when the error leads to an unexpected result. This indicates that coding errors may not only increase the dispersion of results observed in empirical research but may bias the scientific inquiry. The results reinforce the necessity of policies that increase transparency in empirical science.
@@ Appendix A Additional figures and tables
Prior effect 0.130 0.131 0.128 -0.003 0.298 769
@@
Score Q1-Q2 5.127 5.112 5.143 0.031 0.674 805
(1.042) (1.036) (1.050) [0.074]
P-value joint test 0.282
@@
Econometrics 0.002 0.000 -0.002 0.318 1065
[0.002]
Prior effect 0.108 0.097 -0.011 0.558 1065
[0.019]
@@
Effect $0.0138$ $0.0167$ $0.0092$
(s.e.) ($0.0072$) ($0.0076$) ($0.0134$)
[p-value] [$0.0498$] [$0.0228$] [$0.4718$]
{p-value2} {0.02700.0270} {0.02110.0211} {0.22820.2282}
N Obs $805$ $463$ $342$
@@
Table A.4: Heterogeneities - qualified sample
Men Women No Cluster Clustered Below median FS12 Above median FS12
(1) (2) (3) (4) (5) (6)
Effect $0.0098$ $0.0218$ $0.0000$ $0.0142$ $0.0125$ $0.0140$
(s.e.) ($0.0083$) ($0.0134$) - ($0.0075$) ($0.0090$) ($0.0108$)
[p-value] [$0.2183$] [$0.1035$] - [$0.0525$] [$0.1849$] [$0.1423$]
{p-value2} {0.10680.1068} {0.05410.0541} - {0.02720.0272} {0.09410.0941} {0.06900.0690}
N Obs $489$ $309$ $41$ $764$ $403$ $402$
Spotted-1st $0.086$ $0.060$ $0.000$ $0.080$ $0.042$ $0.113$
Spotted-2nd $0.004$ $0.006$ $0.000$ $0.006$ $0.005$ $0.005$
  • Notes: Table presents the results for the combined estimator for different sub-samples of the qualified data. Each column represents the estimates for a different sub-sample, displaying heterogeneity by gender (columns 1 and 2), whether they clustered the standard errors in the screening questions (3 and 4), and below/above the median score in the first two questions (5 and 6). The numbers in parentheses are the robust standard errors. In brackets and curly brackets, the p-values for the bilateral and unilateral tests were obtained using bootstrap with 30,000 replications.
Table A.5: Additional Heterogeneities
Sample Benchmark Master Econometrics
(1) (2) (3)
Effect $0.0138$ $0.0151$ $0.0152$
(s.e.) ($0.0072$) ($0.0076$) ($0.0079$)
[p-value] [$0.0498$] [$0.0449$] [$0.0496$]
{p-value2} {0.02700.0270} {0.02510.0251} {0.02780.0278}
N Obs $805$ $766$ $729$
  • Notes: Table presents the results for the combined estimator for three different samples. The first column presents the results for the benchmark results (qualified sample). The second and third columns restrict the sample for individuals with master’s degree or above and for those that had already taken an econometrics course. We do not present the results for those without a master’s degree or without an econometrics course because these groups are very small. For the qualified sample only 38 individuals do not have a master’s degree and 75 did not take an econometrics course. The numbers in parentheses are the robust standard errors. In brackets and curly brackets, the p-values for the bilateral and unilateral tests were obtained using bootstrap with 30,000 replications.

### v2

@@ 7. Gender
The first three panels split the individuals according to different ways of assessing their coding abilities. Whether they have correctly clustered the standard errors in question 2 (panel 1), their score in the two initial screening questions (panel 2), and their coding score in the partner institute assessment (panel 3). Across the three analysis, the results are very similar. Groups we expect to have better performance (who know how to cluster standard errors and with higher initial scores) have higher probabilities of correctly taking the 99 missing values regardless of the results the coding error would generate. However, these advantages do not make them less vulnerable to the bias. On the contrary, the point estimates are larger for the more trained subgroups although the differences are only significant at 10% for the comparison between those who did or did not correctly clustered the standard errors. The next three panels analyze the results by training and knowledge of econometrics and impact evaluation. We see similar results, those with more training and knowledge (higher impact evaluation scores, with master’s degree, and who have already taken an econometrics course) have larger point estimates, although the differences are not statistically different. In panels 5 and 6, we do not estimate the results for those without master’s degree or who did not took an econometrics course because the sample is so small that we dot enough variation to estimate the combined estimator.
The last panel show heterogeneity by gender, where we observe that women have slightly lower baseline detection probability and larger point estimates for the bias. However, we cannot reject that the results are the same for both samples. In addition to these results, Appendix Table A.3 shows results separately for each wave.
@@ 5 Discussion
Our findings are particularly relevant for placebo tests, where researchers typically expect to find no significant effects—and where expected results are often also seen as favorable. In such cases, if a coding error leads to a statistically insignificant placebo result, researchers may interpret this as confirmation that the test “worked” and may forgo further scrutiny. As a consequence, coding errors may lead to an excess of false-negative findings, masking potential violations of identifying assumptions even when researchers act in good faith.
@@ 6 Conclusion
In recent years, the Economics profession has seen increased concerns about the reproducibility and replicability of research findings. Many journals set up policies requiring data and code availability to increase research transparency. In this paper, we experimentally test whether individuals are more likely to find coding errors when they lead to non-expected results. We find that the probability of spotting a simple and common coding error increases by almost 20% when the error leads to an unexpected result. This indicates that coding errors may not only increase the dispersion of results observed in empirical research but may bias the scientific inquiry. The results reinforce the necessity of policies that increase transparency in empirical science.
@@ Appendix A Additional figures and tables
Score Q1-Q2 5.133 5.121 5.145 0.024 0.744 807
(1.034) (1.019) (1.050) [0.073]
Coding Score 0.150 0.100 0.201 0.101 0.179 733
(1.019) (0.979) (1.057) [0.075]
Impact Evaluation Score 0.158 0.140 0.176 0.036 0.610 751
(0.963) (0.936) (0.992) [0.070]
Prior effect 0.130 0.131 0.128 -0.003 0.229 767
@@
P-value joint test 0.606
@@
Econometrics 0.004 0.000 -0.004 0.157 1036
[0.003]
Coding Score 0.161 0.128 -0.034 0.124 1036
[0.022]
Impact Evaluation Score 0.095 0.073 -0.022 0.197 1036
[0.017]
Prior effect 0.085 0.086 0.001 0.952 1036
[0.017]
@@
Effect $0.0141$ $0.0172$ $0.0093$
(s.e.) ($0.0073$) ($0.0077$) ($0.0135$)
[p-value] [$0.0514$] [$0.0273$] [$0.4652$]
{p-value2} {0.02780.0278} {0.02240.0224} {0.22850.2285}
N Obs $788$ $450$ $338$
@@
Table A.4: Heterogeneities — qualified sample
Effect Std Error P-value N Obs Spotted First Diff P-value
Benchmarks
aEntire Sample 0.0097 (0.0065) [0.1237] 944 0.068 -
aQualified Sample 0.0141 (0.0073) [0.0514] 788 0.078 -
1. Clustered SE?
aYes 0.0145 (0.0077) [0.052] 747 0.081
2. Score in Q1–Q2
aBelow Median 0.0131 (0.0093) [0.1763] 390 0.043 0.9441
aAbove Median 0.0141 (0.0108) [0.1461] 398 0.113
3. Coding Score
aBelow Median 0.0035 (0.0100) [0.7003] 362 0.074 0.4690
aAbove Median 0.014 (0.0105) [0.2005] 362 0.102
4. Impact Evaluation Score
aBelow Median 0.0154 (0.0121) [0.1637] 365 0.087 1.3871
aAbove Median 0.0078 (0.0089) [0.4067] 371 0.063
5. Master’s degree
aYes 0.0154 (0.0077) [0.0454] 750 0.078 -
6. Econometrics course
aYes 0.0156 (0.0081) [0.0476] 714 0.084 -
7. Gender
aMen 0.0101 (0.0085) [0.2128] 477 0.089 0.4519
aWomen 0.0221 (0.0135) [0.1055] 304 0.061
  • Notes: Table presents the results for the combined estimator for different samples. The columns show respectively the point estimate, standard errors (using bootstrap with 30,000 replications) , p-value, number of observations, proportion of individuals spotting the error in the first question (in the control group), and the p-value testing whether the difference of estimates is different than zero. Each row represents the estimates for a different sub-sample. The first two lines are the benchmark results for the entire and qualified samples. Each panel investigates heterogeneity by whether individuals clustered the standard errors in question 2, score in the first two initial questions, coding score, score on the impact evaluation assessment, whether they have a master’s degree, have taken an econometrics course, and by gender. In panels 1, 5, and 6 we do not display results for those who did not cluster, without masters or without econometrics because the sample is very small.